3. Two-factor analysis of variance Emphasis on calculations Aa Aa W. Thomas Boyce, a professor and pediatrician at the University of British Columbia, Vancouver, has studied interactions between individual differences in physiology and differences in experience in determining health and well-being. Dr. Boyce found that some children are more sensitive to their environments. They do exceptionally well when the environment is supportive but are much more likely to have mental and physical health problems when the environment has challenges You decide to do a similar study, conducting a factorial experiment to test the effectiveness of one environmental factor and one physiological factor on a physical health outcome. As the environmental factor, you choose two levels of stressful life events. As the physiological factor, you choose three levels of immune reactivity. The outcome is number of respiratory illnesses in the previous 12 months, and the research participants are kindergartners. You conduct a two-factor ANOVA on the data. The two-factor ANOVA involves several hypothesis tests. Which of the following are null hypotheses that you could use this ANOVA to test? Check all that apply Stressful life events have no effect on number of respiratory illnesses. □ Immune reactivity has no effect on number of respiratory illnesses. The effect of stressful life events on number of respiratory illnesses is no different from the effect of immune reactivity There is no interaction between stressful life events and immune reactivity The results of your study are summarized by the corresponding sample means below. Each cell reports the average number of respiratory illnesses for 9 kindergartners. Factor B: Immune Reactivity Medium Low High M 3.22 T 29 SS= 1.5556 M= 2.44 T 22 SS= 2.2222 M 2.22 T=20 SS= 1.5556 Low TROwi-71 Factor A: X2 415 Stressful Life Events M 2.89 T=26 SS = 0.8889 M= 2.78 T 25 SS = 1.5556 M = 2.78 T=25 SS = 1.5556 High TRow2 = 76 TCOLI = 55 COL2 47 TcoL3 45
media%2Fe35%2Fe35ac63b-cec8-4079-bf8e-c1
media%2F8af%2F8af5ca0b-ae24-4461-b107-25
Answer #1

According to the question, there are two factors, such as immune reactivity and stressful life events. Immune reactivity hasdf. = Number of levels in factor A-1 = 2-1 = 1 dfg = Number of levels in factor B-1 = 3-1 = 2 (Number of levels in factor A -MS AB FB MS-invest 0.9629 0.1944 = 4.95 Therefore, F-ratio of the interaction is 4.95. The complete ANOVA table is shown beloExcel function is used to obtain the critical value for interaction corresponding to 2 numerator degrees of freedom and 48 de